Using Python 3

Implement the calculator for the date of Easter.

The following algorithm computes the date for Easter Sunday for any year between 1900 to 2099.

Ask the user to enter a year. Compute the following:

a = year % 19

b = year % 4

c = year % 7

d = (19 * a + 24) % 30

e = (2 * b + 4 * c + 6 * d + 5) % 7

dateOfEaster = 22 + d + e

Special note: The algorithm can give a date in April. Also, if the year is one of four special years (1954, 1981, 2049, or 2076) then subtract 7 from the date.

Your program should print an error message if the user provides a date that is out of range.
Answer #1
year = int(input('Enter year: '))
if year < 1900 or year > 2099:
    print('Error. Invalid year. Year should be between 1900 and 2099')
else:
    a = year % 19
    b = year % 4
    c = year % 7
    d = (19 * a + 24) % 30
    e = (2 * b + 4 * c + 6 * d + 5) % 7
    dateOfEaster = 22 + d + e
    if year in [1954, 1981, 2049, 2076]:
        dateOfEaster -= 7
    if dateOfEaster <= 31:
        print("Date of easter is March-" + str(dateOfEaster))
    else:
        print("Date of easter is April-" + str(dateOfEaster-31))

\color{blue}Hello,\;Please\;let\;me\;know\;if\;you\;need\;me\;to\;make\;any\;changes. \\ I\;would\;really\;appreciate\;an\;upvote.\;Thank\;you!